OpenStudy (anonymous):

Simplifysquare root of 2 multiplied by the cube root of 2.

OpenStudy (anonymous):

Fan and medal

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@Preetha @christianfilms11

OpenStudy (anonymous):

@dan815 @dtuniyants

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

2 to the power of 1 over 6 2 to the power of 2 over 3 2 to the power of 5 over 6 2 to the power of 7 over 6

OpenStudy (anonymous):

sqrt2=(6root) of 8 cuberoot 2=(6root) 4 (6root 8)*(6root 4)=(6root)(8*4)=(6root)(32)

OpenStudy (anonymous):

so the answer is what

OpenStudy (anonymous):

(6root)32

OpenStudy (anonymous):

\[\sqrt[6]{32}\]

OpenStudy (anonymous):

when i did that i got 33.9

OpenStudy (mertsj):

\[\sqrt{2}\sqrt[3]{2}=2^{\frac{1}{2}}2^{\frac{1}{3}}=2^{\frac{1}{2}+\frac{1}{3}}=2^{\frac{3}{6}+\frac{2}{6}}=2^{\frac{5}{6}}\]